# SurgiTrack: Fine-Grained Multi-Class Multi-Tool Tracking in Surgical Videos

### v1

@@ 1 Introduction
Recently, a new dataset known as CholecTrack20 [22] was introduced providing the multi-class multi-tool tracking requirements. This dataset also formalized three different perspectives of trajectories capturing (1) the life-long intraoperative use of tools, (2) intracorporeal cycle of the tools within the body, and (3) the visibility lifespan of the tools within the camera field of view (FoV) as shown in Fig. 1. Simultaneously tracking tools across these three perspectives is termed multi-perspective tracking. The CholecTrack20 dataset provides rich multi-perspective tracking annotations adaptable to diverse surgical needs, however, to date, no deep learning model has been explored on this dataset for automatic tool tracking.
@@
Inspired by our findings, we propose SurgiTrack, a novel deep learning approach for surgical tool tracking. SurgiTrack casts the tool operators as the approximation of the tools’ originating directions. It then employs an Attention Mechanism to encode tool motion direction, effectively emulating the unseen surgeon operator’s hands or trocar placements for tool re-identification. Our model design allows for an alternative self-supervision of the direction estimator with a comparable performance to its supervised counterpart. This technique ensures that our method can be explored on surgical datasets where the tool operator labels are unavailable. Finally, to account for the multi-perspective nature of tool trajectories, our network associate tracks using a harmonizing bipartite matching graph algorithm, which, aside from the usual linear assignment, resolves identity conflicts across the track perspectives and improves the accuracy of the track ID re-assignment in general.
@@
Fig. 2: Overview of CholecTrack20 dataset showing localization, tracking, and associated labels.
@@ 2 Related Work
Amidst the evolution of tracking methods from rule-based and marker-based approaches [45, 46, 47] to markerless techniques [28, 48] and machine learning methods [49, 15, 9, 7, 26, 10], our work is positioned at the forefront, leveraging deep learning advancements like Siamese networks and attention-based networks [41, 43]. Notably, our focus on direction features, capturing the tool handling direction by surgeon operators, sets our approach apart, offering a more suitable discriminator for refined tool tracking accuracy in scenarios where appearance [50] or similarity [39] features may prove less robust. Contrary to the joint detection and embedding (JDE) setting [32, 31], adopting an online separate detection and tracking (SDE) approach [51, 52, 53, 37, 36, 39], where detection precedes tracking, aligns with the essence of an independently trained tracker that assumes prior detection results. Anchored in dataset like CholecTrack20 [22, 54], which offers both multi-perspective trajectories and MCMOT requirements in endoscopic videos, our research aligns with the ongoing efforts [18, 17, 19, 21] to address the complexities of surgical tool tracking, contributing to the broader landscape of advancements in this critical domain.
@@ 3 Dataset
Our research is conducted on CholecTrack20 [22] dataset consisting of 20 videos of laparoscopic procedures that have been fully annotated with detailed labels for multi-class multi-tool tracking  [^1] Dataset will be released before this paper publication.. The dataset, illustrated in Fig. 2, provides track identities across 3 perspectives of track definition: (1) visibility trajectory of a tool within the camera scope, (2) intracorporeal trajectory of a tool while within a patient’s body, and (3) life long intraoperative trajectory of a tool. It also provides detailed labels for each tool such as spatial bounding box coordinates, class identity, operator identity, phase identity, frame visual conditions such as occlusion, bleeding, and presence of smoke statuses, among others. The annotated tool categories are grasper, bipolar, hook, scissors, clipper, irrigator and specimen bag. The annotated tool operators are main surgeon left hand (MSLH), main surgeon right hand (MSRH), assistant surgeon right hand (ASRH) and null operator (NULL). The annotations are provided at 1 frame per second (FPS) consisting of 35K frames and 65K instance tool labels. Raw videos, recorded at 25 FPS, are provided for inference. The dataset is available on Synapse [54]. More information on the dataset can be found in [22].
@@ 4 Methods
We propose a deep learning method, named SurgiTrack, for surgical tool tracking based on tool direction of motion features. The architecture is conceptually divided into the main components of object tracking: spatial detection and data association, with the later further split into re-identification feature modeling and track identity matching, as illustrated in Fig. 3(a).
@@ 4.1 Spatial Detection
The spatial detection module is responsible for detecting the tools in each frame of a laparoscopic video. We use YOLOv7 [55], a single shot object detector renowned for balancing accuracy and real-time inference, as the backbone for our detection module. This choice emanates from our benchmark study of SOTA detectors on CholecTrack20 dataset (Table 2) where it performs the best.
Pretrained on the Crowdhuman (CH) [56] and MOT20 datasets [57], then finetune on the CholecTrack20 data, YOLOv7 produces tool class identities $C=[0,1,...,6]$ and bounding boxes $B=[X,Y,W,H]$, followed by Non-Maximum Suppression (NMS) at a 0.3 threshold to filter out redundant detections. The predicted bounding boxes are then used to crop out the tool regions from the RGB image for the Re-ID phase.
@@ 4.2.2 Direction Estimation
We model the conditional probability of estimating the direction $\Theta$ of a tool from its bounding box location $B$ following the Bayes rule in Equation 1, implemented using the Attention Mechanism ${\bf A}$ in Equation 2.
$P(\Theta|B)=\frac{P(B|\Theta)\times P(\Theta)}{P(B)}$ (1)
@@
Here, the likelihood $P(B|\Theta)$ of the tool location $B$ given the direction $\Theta$ transforms to the attention alignment function ${\bf KQ^{T}}$ that learns how much attention (weight) each trocar direction should get based on the current tool location. The query ${\bf Q}$ and the key ${\bf K}$ terms are functions that learn the direction vector and the bounding box features respectively. The value ${\bf V}$ models the prior probability $P(\Theta)$ distribution of the directions. And the evidence $P(B)$ which is the probability distribution of the tool locations serves as a scaling term ($\sqrt{d_{\bf K}}$) that normalizes the function. Though the prior probability and evidence can be calculated from the data, for a fully differentiable network, we allow the model to learn these from several iterations on the train data. The ${\bf K,V,Q}$ features are generated from the images via convolutional neural network layers, in this case we use an EfficientNet-b0 [58].
@@ 4.2.3 Operator Estimation
The tool operators $O$ is in turn estimated from the learnt direction $\Theta$ and the category $C$ of the tools. We compute the posterior probability of the operator using Bayes rule in Equation 3 (we omit the normalization term for simplicity without the loss of generality) and model this using a fully connected layer as show in Equation 4.
$P(O|\Theta,C)={P(\Theta|O,C)\times P(O|C)}$ (3)
$O={\bf w}[{\bf A},C]+{\bf b}$ (4)
The likelihood of the direction $P(\Theta|O,C)$ translates to the output of an attention network ${\bf A}$ as in Equation 1. We model the prior probability distribution of the operator per category $P(O|C)$ using the weight ${\bf w}$ of a fully connected layer and add a bias term ${\bf b}$ to complete the implementation. The operator estimation linear layer takes an embedding of the attention-based direction features enhanced by the tool categories $[{\bf A},C]$ as input and produces a vector of logits corresponding to the probability of each operator given a tool.
@@ 4.3 Identity Association using Multi-Perspective Matching
The role of the identity association module is to link tool detections across frames, creating coherent tracks for surgical tools. It is a critical step in understanding how these tools move and interact during procedures. To tackle this challenge effectively, we introduce a novel approach called Harmonizing Bipartite Graph Matching (HBGM). This method is designed to handle the diverse perspectives of surgical tool trajectories without conflicts, ensuring that every tool is tracked accurately throughout the surgery. HBGM builds upon the Bipartite graph matching (BGM) algorithm [59], which is a powerful tool in its own right. What sets HBGM apart is its ability to manage multiple track states for each tool instance. These states include track visibility, whether a tool is currently in use within or out of the camera’s view, within or out of the body. The stages of HBGM are illustrated in Figure 3(c) and discussed as follows:
@@ 4.3.1 Pairwise Matching
First, we calculate pairwise distances between predicted motion directions and detection boxes. A bipartite matching algorithm is used for linking detections based on motion direction, refining with spatial distance and tool category information. A second bipartite association is optionally performed using low-confidence detections, as proposed in [38].
@@ 4.3.2 Track Harmonization
The magic of HBGM happens during the harmonization stage, where it skillfully handles track ID assignment. It identifies, from its rich history, when tools move out of the camera’s view (OOCV), temporally leave the surgical field (out-of-body OOB), or are no longer in use by maintaining a unique identity per temporal perspective of a tool trajectory and enforcing referential relation across perspectives. The linear matching employs higher threshold for re-identifying OOCV tools and lower threshold for OOB tools. By considering these different trajectory perspectives, HBGM ensures there are no conflicts in the track identities.

### v3

@@ 1 Introduction
Article to appear in Medical Image Analysis Journal, 2025
@@
Recently, a new dataset known as CholecTrack20 [36] was introduced providing the multi-class multi-tool tracking requirements. This dataset also formalized three different perspectives of trajectories capturing (1) the life-long intraoperative use of tools, (2) intracorporeal cycle of the tools within the body, and (3) the visibility lifespan of the tools within the camera field of view (FoV) as shown in Fig. 1. Simultaneously tracking tools across these three perspectives is termed multi-perspective tracking. The CholecTrack20 dataset provides rich multi-perspective tracking annotations adaptable to diverse surgical needs, however, to date, no deep learning model has been explored on this dataset for automatic surgical tool tracking.
@@
Inspired by our findings, we propose SurgiTrack, a novel deep learning approach for surgical tool tracking. SurgiTrack casts the tool operators as the approximation of the tools’ originating directions. It then employs an Attention Mechanism to encode tool motion direction, effectively emulating the unseen surgeon operator’s hands or trocar placements for tool re-identification. Our model design allows for an alternative self-supervision of the direction estimator with a comparable performance to its supervised counterpart. This technique ensures that our method can be explored on datasets where the tool operator labels are unavailable. Finally, to account for the multi-perspective nature of tool trajectories, our network associate tracks using a harmonizing bipartite matching graph algorithm, which, aside from the usual linear assignment, resolves identity conflicts across the track perspectives and improves the accuracy of the track ID re-assignment in general.
@@
Fig. 2: Overview of CholecTrack20 dataset showing localization, tracking, and associated labels [36].
Table 1: Summary of the CholecTrack20 dataset statistics [36].
Details Average track length / number of tracks per perspective Surgical visual challenges
Params Counts Tools Intraoperative Intracorporeal Visibility Events Counts
Videos 20 Grasper 582.0 / 59 209.4 / 164 25.7 / 1334 Occlusion 22958
Total duration 14h 3m 1s Bipolar 118.0 / 20 30.6 / 78 24.6 / 96 Bleeding 18698
Tool categories 7 Hook 1063.7 / 20 181.8 / 118 87.5 / 244 Smoke 2388
Phase categories 7 Scissors 44.6 / 20 27.0 / 34 23.4 / 46 Crowded 6737
Operator categories 4 Clipper 71.7 / 20 36.8 / 40 31.9 / 46 Blurring 299
Labeled frames (1 FPS) 35K Irrigator 90.8 / 20 25.9 / 56 13.4 / 108 Light reflection 74
Total frames (25 FPS) 1.3M Spec.Bag 175.2 / 20 166.8 / 20 44.9 / 78 Fouled lens 2196
Bounding boxes 65.2K Total 2146.0 / 179 678.3 / 510 251.4 / 1952 Poor camera coverage 357
@@ 2 Related Work
Amidst the evolution of tracking methods from rule-based and marker-based approaches [31, 24, 9] to markerless techniques [41, 61] and machine learning methods [44, 66, 2, 40, 50, 15], our work is positioned at the forefront, leveraging deep learning advancements like Siamese networks and attention-based networks [51, 34]. Notably, our focus on direction features, capturing the tool handling direction by surgeon operators, sets our approach apart, offering a more suitable discriminator for refined tool tracking accuracy in scenarios where appearance [42] or similarity [56] features may prove less robust. Contrary to the joint detection and embedding (JDE) setting [64, 57], adopting an online separate detection and tracking (SDE) approach [7, 32, 19, 1, 18, 56], where detection precedes tracking, aligns with the essence of an independently trained tracker that assumes prior detection results. Anchored in dataset like CholecTrack20 [36], which offers both multi-perspective trajectories and MCMOT requirements in endoscopic videos, our research aligns with the ongoing efforts [35, 46, 21, 47] to address the complexities of surgical tool tracking, contributing to the broader landscape of advancements in this critical domain.
@@ 3 Dataset
Our research is conducted on CholecTrack20 [36] dataset consisting of 20 videos of laparoscopic procedures that have been fully annotated with detailed labels for multi-class multi-tool tracking. The dataset, illustrated in Fig. 2, provides track identities across 3 perspectives of track definition: (1) visibility trajectory of a tool within the camera scope, (2) intracorporeal trajectory of a tool while within a patient’s body, and (3) life long intraoperative trajectory of a tool. Intraoperative tracking not only re-identifies tools out of camera view (OOCV) as done in intracorporeal tracking but also maintains their trajectory when out of body (OOB). In the CholecTrack20 dataset, OOB is detected/annotated either by visually observing the tool exit the trocar, inferring from another tool entering through the same trocar, or noting that the initial tool releases its grasp while out of camera focus. The dataset also provides detailed labels for each tool such as spatial bounding box coordinates, class identity, operator identity, phase identity, frame visual conditions such as occlusion, bleeding, and presence of smoke statuses, among others. The annotated tool categories are grasper, bipolar, hook, scissors, clipper, irrigator and specimen bag. The annotated tool operators are main surgeon left hand (MSLH), main surgeon right hand (MSRH), assistant surgeon right hand (ASRH) and null operator (NULL). The annotations are provided at 1 frame per second (FPS) consisting of 35K frames and 65K instance tool labels. Raw videos, recorded at 25 FPS, are provided for inference. The dataset is available on Synapse [37]. A summary of the CholecTrack20 dataset statistics showing the size of key elements in the dataset, number of challenging scenarios, average track duration and number of tracks for each tool category per perspective is presented in Table 1. More information on the dataset can be found in Nwoye et al. [36].
@@ 4 Methods
We present SurgiTrack, a deep learning method for surgical tool tracking based on tool direction of motion features. SurgiTrack is designed as a multi-class multi-object tracking (MCMOT) model capable of tracking tools jointly across multiple trajectory perspectives, namely, visibility, intracorporeal, and intraoperative. The motivation to track beyond camera’s field of view is to offer more flexible trajectories that ensure continuous and reliable identification of surgical tools, tailored to the complex dynamics of a surgical scene, preventing errors and maintaining safety even when tools temporarily move out of view.
The architecture of our proposed tracking model is conceptually divided into the main components of object tracking: spatial detection and data association, with the later further split into re-identification feature modeling and track identity matching, as illustrated in Fig. 3(a).
4.1 Multi-Perspective Tool Track Formalization
Given a video dataset $D=\{S_{1},S_{2},\ldots,S_{N}\}$, where each sequence $S_{i}=\{f_{1},f_{2},\ldots,f_{T}\}$ consists of frames $f_{t}$ containing tools represented by bounding box locations $B_{t}=[B_{1}^{t},B_{2}^{t},\ldots,B_{M}^{t}]$ and associated classes $C_{t}=[C_{1}^{t},C_{2}^{t},\ldots,C_{M}^{t}]$, the objective is to accurately track tool identities over time across multiple perspectives. For each perspective $p\in\{1,2,3\}$, we define an association matrix $A^{(p)}(t)$, where $A^{(p)}_{i,j}(t)=1$ indicates that the $i$-th detected tool in frame $f_{t}$ is associated with the $j$-th detected tool in frame $f_{t+1}$, and $A^{(p)}_{i,j}(t)=0$ otherwise. The final tracking solution involves harmonizing these perspective-specific matrices $\{A^{(1)}(t),A^{(2)}(t),A^{(3)}(t)\}$ to ensure consistent and accurate tool tracking.
@@ 4.2 Spatial Detection
The spatial detection module is responsible for detecting the tools in each frame of a laparoscopic video. The detection output $D_{t}=(B_{i}^{t},C_{i}^{t})$ of frame $f_{t}$ is a pair of bounding box coordinates and class identities, with $B_{i}^{t}=[X_{i}^{t},Y_{i}^{t},W_{i}^{t},H_{i}^{t}]$ containing the center coordinates $(X_{i}^{t},Y_{i}^{t})$ and the spatial sizes $(W_{i}^{t},H_{i}^{t})$ of the boxes and $C_{i}^{t}\in\{0,1,\ldots,6\}$ represents the class identity of the tools.
We use YOLOv7 [54], a single shot object detector renowned for balancing accuracy and real-time inference, as the backbone for our detection module. Pretrained on the Crowdhuman (CH) [48] and MOT20 datasets [14], then finetune on the CholecTrack20 data, YOLOv7 produces tool class identities $C=[0,1,\ldots,6]$ and bounding boxes $B=[X,Y,W,H]$, followed by Non-Maximum Suppression (NMS) at a 0.3 threshold to filter out redundant detections. The predicted bounding boxes are then used to crop out the tool regions from the RGB image for the Re-ID phase.
@@ 4.3.2 Direction Estimation
We model the estimation of the direction $\theta$ of a surgical tool, given its bounding box location $B$, using a Bayesian-inspired framework. Specifically, the relationship between $\theta$ and $B$ follows Bayes’ rule as shown in Equation 1:
$P(\theta|B)=\frac{P(B|\theta)\cdot P(\theta)}{P(B)}$ (1)
Here, $P(B|\theta)$ represents the likelihood of the tool’s bounding box given its direction, $P(\theta)$ is the prior probability of the direction, and $P(B)$ is the evidence, representing the marginal probability of the bounding box across all possible directions.
We approximate this framework using the Attention Mechanism ${\bf A}$ in Equation 2, which captures dependencies between bounding box features and direction:
@@
The Attention Mechanism is a common approach in neural networks for learning dependencies between variables [3]. In this formulation, the likelihood $P(B|\theta)$ is approximated by the attention alignment scores $\text{\it SoftMax}~{}({\bf KQ}^{T}/\sqrt{d_{\bf K}})$, which measure the relevance of each direction based on the bounding box features, i.e., how much attention (weight) each trocar direction should get based on the current tool location. The prior probability $P(\theta)$ is represented by the value vector ${\bf V}$, which learns a distribution over possible directions independent of the bounding box features. The evidence $P(B)$ serves as a normalization term, implicitly handled through the SoftMax operation. The query ${\bf Q}$, key ${\bf K}$, and value ${\bf V}$ are learned embeddings that encode direction vectors, bounding box features, and direction priors, respectively, from the input images using convolutional neural network (CNN) layers. Specifically, we use EfficientNet-b0 [53] to encode these features.
The Attention Mechanism ${\bf A}$ effectively approximates $P(\theta|B)$ by learning the alignment between bounding box features and directional priors within a differentiable framework. Unlike classical Bayesian methods where priors and likelihoods are explicitly defined, this approach allows the network to learn these distributions directly from the data, providing flexibility for modeling complex spatial dependencies. By grounding the formulation in Bayes’ rule and leveraging the interpretability of attention mechanisms, we provide a principled approach to estimating tool direction while maintaining compatibility with modern deep learning architectures.
@@ 4.3.3 Operator Estimation
The tool operator $O$ is estimated as the conditional probability $P(O\mid[{\bf A},C])$, where ${\bf A}$ represents the attention-based direction features (derived from Equation 2), and $C$ is the tool category. This estimation is modeled using a fully connected layer, with the linear transformation defined as:
$y=\mathbf{w}\cdot[{\bf A},C]+\mathbf{b}$ (3)
where, $\mathbf{w}$ and $\mathbf{b}$ are the weight and bias parameters, and the input is the concatenation of ${\bf A}$ and $C$. To obtain the final probability distribution, a Softmax function is applied to $y$:
$P(O\mid[{\bf A},C])=\text{Softmax}(y)$ (4)
This process generates the probability of each operator $O$ given the tool direction and category, enabling operator classification.
@@ 4.4 Identity Association using Multi-Perspective Matching
The role of the identity association module is to link tool detections across frames, creating coherent tracks for surgical tools. It is a critical step in understanding how these tools move and interact during procedures. To tackle this challenge effectively, we introduce a novel approach called Harmonizing Bipartite Graph Matching (HBGM). This method is designed to handle the diverse perspectives of surgical tool trajectories without conflicts, ensuring that every tool is tracked accurately throughout the surgery. HBGM builds upon the Bipartite graph matching (BGM) algorithm [27], which is a powerful tool in its own right. What sets HBGM apart is its ability to manage multiple track states for each tool instance. These states include:
  1. 1.
    New: Assigned at track initialization.
  2. 2.
    Active: Updated when a detection matches a track.
  3. 3.
    Lost: Assigned to an active track with no matching detection at current time. The “Lost” state has a 5-second time-threshold within which the track can be re-activated for all perspectives.
  4. 4.
    Out of Camera View (OOCV): Assigned if a track remains “Lost” for more than the time-threshold. This state marks the end of a visibility trajectory. An “OOCV” track can only be re-activated for intraoperative and intracorporeal perspectives.
  5. 5.
    Out of Body (OOB): Assigned if a tool track is judged to have left the abdomen. This is determined when a track fails a class-direction consistency check (described in Section 4.4.2). The OOB state marks the end of an intracorporeal trajectory. Such track can only be re-activated for intraoperative perspective.
  6. 6.
    Removed: Assigned if a track is initialized erroneously (e.g., after the “New” state, it never progresses to “Active” for a prolonged time) or not reassigned for a long period. Removed tracks are excluded from history against a potential match.
The stages of HBGM are illustrated in Figure 3(c) and discussed as follows:
@@ 4.4.1 Pairwise Matching
To match tools across frames, we use pairwise distances between predicted motion directions ($\theta$) and detection boxes ($B$), while ensuring class ($C$) consistency. For each predicted tool $p_{i}^{t}=\{B_{i}^{t},C_{i}^{t},\theta_{i}^{t}\}$ in frame $f_{t}$, and each active track $q_{j}^{t-1}=\{B_{j}^{t-1},C_{j}^{t-1},\theta_{j}^{t-1},~{}\text{IDs}\}$ in the history trajectories up to frame $f_{t-1}$, the distance metric $Dist_{i,j}$ is defined as:
$\text{Dist}_{i,j}=\alpha\cdot\text{Dist}_{\text{spatial}}(B_{i}^{t},B_{j}^{t+1})~{}+\\ \beta\cdot\text{Dist}_{\text{direction}}(\theta_{i}^{t},\theta_{j}^{t+1})~{}+\\ \gamma\cdot\text{Dist}_{\text{class}}(C_{i}^{t},C_{j}^{t+1}),$ (7)
where
$\text{Dist}_{\text{spatial}}(B_{i}^{t},B_{j}^{t+1})=1-\frac{|B_{i}^{t}\cap B_{j}^{t+1}|}{|B_{i}^{t}\cup B_{j}^{t+1}|},$ (8)
is the Intersection over Union (IoU) similarity between the bounding boxes $B_{i}^{t}$ and $B_{j}^{t+1}$. Alternatively, $Dist_{spatial}(B_{i}^{t},B_{j}^{t+1})$ can be computed as the Euclidean distance between bounding box centers: $\sqrt{(X_{i}^{t}-X_{j}^{t+1})^{2}+(Y_{i}^{t}-Y_{j}^{t+1})^{2}}$,
$\text{Dist}_{\text{direction}}(\theta_{i}^{t},\theta_{j}^{t+1})=1-\frac{\theta_{i}^{t}\cdot\theta_{j}^{t+1}}{\|\theta_{i}^{t}\|\cdot\|\theta_{j}^{t+1}\|},$ (9)
is the angular difference between motion directions, and
$\text{Dist}_{\text{class}}(C_{i}^{t},C_{j}^{t+1})=\begin{cases}1-{S}&\text{if }C_{i}^{t}=C_{j}^{t+1}\\ \tau&\text{if }C_{i}^{t}\neq C_{j}^{t+1},\end{cases}$ (10)
ensures that tools of different classes are not matched ($S$ is the class probability score of the tool and $\tau$ is set above matching threshold). Here, $\alpha$, $\beta$, and $\gamma$ are weights that balance the influence of spatial, motion, and class distances, respectively. A bipartite matching algorithm is then used to minimize $\text{Dist}_{i,j}$ and establish correspondences. Optionally, a second bipartite association is performed for low-confidence detections, following the method proposed by Zhang et al. [63].
@@ 4.4.2 Track Harmonization
The magic of HBGM happens during the harmonization stage, where it skillfully handles track ID assignment ensuring consistent tool tracking across multiple perspectives — namely intraoperative, intracorporeal, and visibility. The harmonization process relies on a structured approach involving (a) class-direction consistency checks, (b) hierarchical ID assignment, and (c) referential integrity to effectively manage tool trajectories under different states and handle potential conflicts in track identification and state updates.
A. Class-Direction Consistency Check: HBGM validates that a track maintains a joint matching of class ($C$) and direction ($\theta$) features across time to remain active. This check is specifically for managing the intracorporeal trajectory. A change in the class feature (or class ID) while the direction feature remains unchanged indicates that a new tool $j$ has been introduced through the same trocar as previous tool $i$, marking the end of the tool $i$ intracorporeal trajectory and the start of a new one:
$\text{If }\theta_{i}^{t}=\theta_{j}^{t-1}\text{ and }C_{i}^{t}\neq C_{i}^{t-1},\text{ then }\text{State}_{j}^{t-1}=\text{OOB}.$ (11)
This approach enforces a one-to-one correspondence between track direction and class features for more stable and accurate tracking.
B. Hierarchical ID Initialization and State Updates: HBGM employs a hierarchical approach to manage new ID assignments and state updates effectively across perspectives. At the topmost level is the intraoperative trajectory ($p=3$) followed by the intracorporeal trajectory ($p=2$), and lastly the visibility trajectory ($p=1$).
  1. 1.
    ID Initialization: A new ID in a higher perspective ($p=3$) triggers new IDs in all lower perspectives ($p=2$, $p=1$). Similarly, new IDs initialized in $p=2$ are also initialized in $p=1$.
  2. 2.
    State Updates: State changes in lower perspectives are propagated in higher ones. For example,“Lost” in $p=1$ implies “Lost” in $p=2$ and $p=3$. “OOCV” in $p=2$ implies “OOCV” in $p=3$. Hierarchical state update is excluded for “New” state, which is treated as ID initialization.
C. Referential Integrity: Once a tool’s track is established in one perspective, every changes in its state is concurrently reflected across all other perspectives. HBGM improves consistent track assignment for unmatched detections and untracked tracks by cross-referencing of historical data across perspectives. This minimizes ID switches and facilitates track recovery after occlusion, OOCV or OOB. Also, persistent trajectory patterns in $p=3$ is relied on to dynamically adjust thresholds for the inactive tracks, aiding in robust recovery.